Calculate what $750 would grow to at 6% per year compounded daily for 9 years. (Use the Table 12.2.) (Round your final answer to the nearest cent.)
Amount:

Found 2 solutions by ikleyn, Theo:
Answer by ikleyn(52908)

.

I have no Table 12.2 you refer to.

And no one tutor at this forum has no it . . .



Answer by Theo(13342)

not sure what table you are looking at.

however, the formula is:

f = p * (1 + r) ^ n
f is the future value
p is the present value
r is he interest rate per time period
n is the number of time periods.

normally, 365 days per year is assumed; the interest rate per year is used, not the percent.
therefore, the formula becomes:

f = 750 * (1 + .06/365) ^ (9 * 365) = 1286.948032.
round to the nearest penny to get 1286.95.

the interest rate per time period turns out to be .06/365 equals:
.0001643835616.

the number of time periods turns out to be 9 * 365 = 3285.

your table might combine the two and tell you that 9 years at 6% per year compounded daily = (1 + .001643835616 ^ 3285 = 1.715930709.

that would be the factor that you multiply the present value of 750 by.

your table might also be rounded to 3 or 4 decimal places.

note that 750 * (1 + .06/365) ^ (9 * 365) = 750 * 1.715930709 = 1286.948032.

there might be some slight differences due to rounding of intermediate results.

for example, the table might show 1.7159.
in that case, your answer might be 750 * 1.7159 = 1286.93 rounded to the nearest penny.